ANALYSIS OF ROTATION-VIBRATION SPECTRA OF SMALL MOLECULES EMBEDDED IN PARAHY-DROGEN CRYSTALS

\underline{HIROMICHI HOSHINA}, MIZUHO FUSHITANI and TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, JAPAN.

We have been using solid parahydrogen (p-H_{2}) as a matrix for high-resolution infrared spectroscopy. Small molecules embedded in solid p-H_{2} are found to exhibit completely free quantized rotational motion. In previous works, we have fully analyzed the high-resolution rotation-vibration spectra of spherical top molecules (CH_{4} and CD_{4}) in solid p-H_{2}.^{a}^{b} We found that the rotational constants of these molecules decreases 10%∼20% from the gas phase values due to the interaction between the guest molecules and surrounding hydrogen molecules. This paper will focus on the analysis of the spectra of symmetric top molecules such as CD_{3} and CD_{3}H. The ν_{3} rotation-vibration transition of CD_{3} shows complic! ated spectral structure due to the quantized rotational motion and the crystal field splittings. The analysis of the spectrum reveals that the reduction of the rotational constant C is three times larger than that of B. The origin of the reduction will be discussed.

aT. Momose, M. Miki, T. Wakabayashi, T. Shida, M.-C. Chan, S. S. Lee, T. Oka J. Chem. Phys. 107 7707 (1997)

bH. Hoshina, T. Wakabayashi, T. Momose, and T. Shida, J. Chem. Phys. 110 5728 (1999)